# Test-time Adaptive Hierarchical Co-enhanced Denoising Network for Reliable Multimodal Classification

### v1

@@ Abstract
Reliable learning on low-quality multimodal data is a widely concerning issue, especially in safety-critical applications. However, multimodal noise poses a major challenge in this domain and leads existing methods to suffer from two key limitations. First, they struggle to reliably remove heterogeneous data noise, hindering robust multimodal representation learning. Second, they exhibit limited adaptability and generalization when encountering previously unseen noise. To address these issues, we propose Test-time Adaptive Hierarchical Co-enhanced Denoising Network (TAHCD). On one hand, TAHCD introduces the Adaptive Stable Subspace Alignment and Sample-Adaptive Confidence Alignment to reliably remove heterogeneous noise. They account for noise at both global and instance levels and enable jointly removal of modality-specific and cross-modality noise, achieving robust learning. On the other hand, TAHCD introduces test-time cooperative enhancement, which adaptively updates the model in response to input noise in a label-free manner, improving adaptability and generalization. This is achieved by collaboratively enhancing the joint removal process of modality-specific and cross-modality noise across global and instance levels according to sample noise. Experiments on multiple benchmarks demonstrate that the proposed method achieves superior classification performance, robustness, and generalization compared with state-of-the-art reliable multimodal learning approaches.
I Introduction
Humans perceive and understand the world through multiple sensory channels, such as vision, hearing, and smell. Inspired by this capability, multimodal learning [29, 48, 27, 2] has gained substantial attention in deep learning. With recent advances in sensing, video technologies, and data processing, large volumes of multimodal data have become available, further accelerating progress in this field. By effectively fusing complementary information across modalities, multimodal learning enables more comprehensive and discriminative representations. As a result, it significantly improves performance across a wide range of tasks, including disease diagnosis [12, 52], image–text classification [47, 3, 7], video classification [26, 19, 23], and emotion analysis [37, 36, 33].
@@ I Introduction
While multiple modalities enrich representation learning, they inevitably introduce more complex noise, typically arising from sensor errors, environmental interference, and data collection errors. In real-world scenarios, multimodal noise exhibits two key characteristics: (1) Heterogeneity. Previous work [46] identifies two main types of multimodal noise: modality-specific noise and cross-modality noise. Modality-specific noise arises independently in each modality from sensor errors or environmental disturbances, while cross-modality noise stems from low-quality data due to weakly aligned or misaligned multimodal samples [45, 5]. (2) Unpredictability. Noise often evolves over time as new samples are incrementally acquired, potentially introducing previously unseen noise. For example, in the medical domain, unexpected device failures may generate unfamiliar noise in newly collected patient data. In autonomous driving, sensors often face varying weather and urban conditions. In Internet data collection, diverse image–text–mismatched posts continuously emerge over time.
@@
In the presence of complex multimodal noise, conventional multimodal methods struggle to produce reliable results, making them unsuitable for safety-critical applications such as computer-aided disease diagnosis and autonomous driving [9, 31, 50]. Accordingly, many methods improve the reliability of multimodal learning by removing noise and uninformative feature in data. Among them, [13, 14, 11, 12, 50, 52, 47, 3] address modality-specific noise via confidence-based feature learning and fusion. However, they generally assume well-aligned inputs and overlook the presence of cross-modality noise. Conversely, [17, 24, 32, 44, 43] eliminate cross-modality noise via alignment and rectification, but their lack of modality-confidence awareness leads to limited robustness under modality-specific noise. Therefore, these methods generally face the following limitations. (1) Insufficient robustness to heterogeneous noise: They struggle to achieve joint removal of modality-specific and cross-modality noise. (2) Limited generalization: they can only learn and update on limited training data, making it difficult to handle previously unseen noise.
@@
To address these limitations, we propose Test-time Adaptive Hierarchical Co-enhanced Denoising Network (TAHCD). TAHCD first removes heterogeneous noise by considering it at both global and instance levels. As shown in Fig. 1, data often contain global-level noise with consistent patterns across samples, such as persistent sensor faults or repeated sampling under identical conditions, and instance-level noise affecting individual samples, e.g., sudden sensor failures or human errors. Accordingly, TAHCD proposes Adaptive Stable Subspace Alignment (ASSA) and Sample-Adaptive Confidence Alignment (SACA) to jointly eliminate modality-specific and cross-modality noise at the global and instance levels, respectively, achieving robust representation learning. Moreover, TAHCD incorporates test-time cooperative enhancement (TTCE) to collaboratively refine global- and instance-level denoising based on input noise in a label-free manner, enabling adaptation to previously unseen noise and substantially improving generalization.
@@
Specifically, ASSA first computes each modality’s feature covariance and performs singular value decomposition. A learnable mask conditioned on the singular values then selects eigenvectors to construct stable feature subspaces, guided by inter-class orthogonality and subspace projection alignment. These constraints encourage the mask to discard directions carrying modality-specific and cross-modality noise across all samples. By aligning modality projections in the constructed stable subspaces, the subspace projection alignment in ASSA improves the reliability of cross-modality noise removal by preventing the model from erroneously aligning uninformative semantic content induced by modality-specific noise. Subsequently, SACA uses the globally denoised features from ASSA to estimate priors that guide the experts in removing heterogeneous noise at the instance level via a confidence-aware asymmetric slack alignment. By reweighting gradients based on modality confidence, this strategy encourages low-confidence modalities to align towards high-confidence ones, enabling reliable cross-modality noise removal under instance-level modality-specific noise. It also adopts a slack alignment scheme that constrains cross-modality discrepancies within a reasonable range learned from the prior, reflecting inherent differences in informative content across modalities. Compared with existing methods that strictly enforce cross-modality consistency, this prevents over-suppressing modality-specific complementary information while removing noise, resulting in more robust multimodal representations.
@@
Building upon the above framework for heterogeneous noise removal, we further propose TTCE that adapts the denoising framework to unseen noise in a label-free manner. Since the instance level can capture noise variations more sensitively than the global level, it can also perceive unseen noise more accurately. Based on this consideration, TTCE first incorporates the instance-level modality-specific and cross-modality noise learned by SACA into the reconstruction of globally denoised modality features produced by ASSA. This enables global-level modality representation learning in ASSA to explicitly leverage more accurate and finer-grained unseen-noise information captured at the instance level, thereby improving global-level unseen-noise removal. The enhanced global-denoised features in ASSA can provide more reliable priors, which in turn improve instance-level unseen-noise learning and removal in SACA. Through multiple iterations of the above ASSA–SACA co-enhancement process, the model achieves progressive update and adaptation to unseen noise.
@@
  • The proposed TTCE in TAHCD cooperatively enhances ASSA and SACA driven by input sample noise in a label-free manner. This enables adaptive improvement of the denoising process at test time for previously unseen noise, thereby enhancing model generalization ability.
II Related Works
@@
II-A Multimodal Learning
The advancement of sensor technology and the development of data transmission methods have led to the emergence of a vast amount of data from diverse sources, types, and formats. This has driven the rapid growth of multimodal learning [35, 2, 51], which has substantially improved machine learning performance across a wide range of tasks and applications [15, 16, 25]. For instance, in biomedical applications, DeepIMV [20] and MOGONET [40] have achieved more precise pathological classification by integrating multiple omics data. In computer vision, CEN [42] has enhanced dense image prediction by fusing multimodal data and adaptively exchanging channels of different modality models. Zhang et al. [48] have guided image outpainting more effectively by balancing the fusion of image and text modalities. In recent years, growing attention has been directed toward addressing the challenges of multimodal learning in real-world scenarios, one of which is the need to handle low-quality multimodal data. [46, 47]. However, conventional multimodal learning methods struggle to produce robust and generalizable outputs when faced with low-quality data, making it difficult to meet the reliability requirements of the model, especially in safety-critical applications. Consequently, we propose a method for reliable multimodal learning that learns robust representations under heterogeneous noise while effectively adapting and generalizing under unseen noise.
II-B Reliable Learning from Low-Quality Multimodal Data
@@
In real-world scenarios, multimodal learning often suffers from low-quality data, which undermines reliability. A major source of such low-quality data is multimodal noise, typically arising from factors such as sensor failures, environmental influences, or human errors. Existing studies [46] categorize multimodal noise into two types, namely modality-specific noise and cross-modality noise, and have developed various methods to address these noise types for reliable multimodal learning.
Many studies mitigate noise in each modality and improved robustness through confidence-based representation learning and fusion methods. For instance, Federici et al. [8] have proposed the multi-view information bottleneck, which has improved the robustness of multi-view learning by retaining information shared by each view. Han et al. [13] have parameterized the evidence of different modality features using Dirichlet distribution and fused each modality at the evidence level using Dempster-Shafer theory. Geng et al. [11] have proposed the DUA-Nets, which achieved uncertainty-based multimodal representation learning through reconstruction. Han et al. [12] have modeled informativeness at the feature and modality levels, achieving trustworthy multimodal feature fusion. Zhang et al. [47] have achieved more robust multimodal fusion by dynamically assigning weights to each modality based on uncertainty estimation. Zheng et al. [49] have achieved trustworthy multimodal classification via integrating feature and label-level confidence. Zou et al. [52] have proposed a novel dynamic poly-attention Network that integrated global structural information for trustworthy multimodal classification. Zhou et al. [50] have introduced a trustworthy multi-view classification framework by enhancing multi-view encoding and confidence-aware fusion. Cao et al. [3] have proposed the Predictive Dynamic Fusion method, which provides theoretical guarantees for reducing the upper bound of generalization error in dynamic multimodal fusion.
Additionally, many studies focus on cross-modality noise removal by modeling alignment and correspondence between modalities. Radenovic et al. [34], Gadre et al. [10], and Sharma et al. [38] have proposed rule-based methods to remove misaligned or weakly aligned data. NCR [18], ALBEF [22], BLIP [21], and SPS [43] have adopt model-based rectification methods, which design models to filter or correct misaligned samples. Additionally, NLIP [17], OSCAR [24], and SMILE [44] have proposed noise-robust regularization methods to mitigate the impact of misaligned samples on the model.
Although the aforementioned methods effectively remove multimodal noise and enhance the reliability of representation learning, they generally exhibit the following limitations. First, they struggle to jointly remove modality-specific and cross-modality noise, resulting in representations with limited robustness under heterogeneous data noise. Moreover, they can only learn and update on limited training data, making them poorly adaptive to unseen noise and limiting generalization. To address these issues, we propose Test-time Adaptive Hierarchical Co-enhanced Denoising Network (TAHCD), which jointly removes modality-specific and cross-modality noise at both global and instance levels, and enables label-free, adaptive updates to model inputs at test time, significantly improving robustness and generalization.
@@
III Proposed Approach
III-A Overview of TAHCD
As illustrated in Fig. 2, TAHCD first employs modality-specific feature encoders to extract representations from each modality. It then introduces Adaptive Stable Subspace Alignment (ASSA) to remove modality-specific and cross-modality noise exhibiting global distributional characteristics. In addition, Sample-Adaptive Confidence Alignment (SACA) is proposed to eliminate modality-specific and cross-modality noise at the instance level. Building upon these components, we further design test-time cooperative enhancement (TTCE) mechanism. It achieves cooperative enhancement between global- and instance-level denoising, driven by sample noise in a label-free manner, improving adaptability and generalization to unseen noise.
III-B Adaptive Stable Subspace Alignment
@@
III-B1 Adaptive Stable Subspace Construction
To facilitate global-level multimodal noise removal, we adaptively eliminate statistically unstable and uninformative directions in the feature distribution of each modality and suppress noise-dominated variations across samples. Specifically, for the $m$-th modality, the input $x^{m}=\{x_{1}^{m},\dots,x_{N}^{m}\}$ is first encoded into the latent space $z^{m}=\{z_{1}^{m},\dots,z_{N}^{m}\}$ by the corresponding modality encoder $\phi_{x}^{m}$: $z^{m}=\phi_{x}^{m}(x^{m})$. Then, the covariance matrix $\Sigma^{m}_{z}$ among features is computed from $z^{m}$, which can be formulated as follows:
@@ III-B1 Adaptive Stable Subspace Construction
$\phi_{\lambda}^{m}(\cdot)$ is a learnable encoder, $\sigma(\cdot)$ denotes the sigmoid function. $h^{m}=\{h_{1}^{m},\dots,h_{N}^{m}\}$ denote the globally denoised representation obtained from the latent feature $z^{m}$ of modality $m$. As shown in Eq. (5), the feature vector $z^{m}$ is first projected onto the covariance principal axes using the orthogonal basis $U_{z}^{m}$. It is then filtered by the learned mask $w^{m}_{\lambda}$ and finally mapped back to the original feature space via $(U^{m}_{z})^{\top}$.
III-B2 Constraints
To ensure that the adaptive stable subspace construction sufficiently eliminates modality-specific and cross-modality noise, we introduce inter-class orthogonality constraint $\mathcal{L}_{\mathrm{o}}$ and subspace projection alignment constraint $\mathcal{L}_{\mathrm{a}}$. The inter-class orthogonality constraint $\mathcal{L}_{\mathrm{o}}$ eliminates modality-specific noise by removing the shared spurious patterns it induces across classes within the same modality, while preserving class-discriminative features, which can be formulated as:
@@ III-B2 Constraints
III-C Sample-Adaptive Confidence Alignment
To provide reliable supervision signals for instance-level noise removal, we first estimate modality-wise feature distributions and cross-modality discrepancy distributions using the globally denoised representations obtained from ASSA. These distributions are then serve as priors to guide sample-adaptive modality-specific and cross-modality noise experts with confidence-aware asymmetric slack alignment for instance-level noise removal.
III-C1 Prior Estimation

### v2

@@ Abstract
Reliable learning of multimodal data (e.g., multi-omics) is a widely concerning issue, especially in safety-critical applications such as medical diagnosis. However, low-quality data induced by multimodal noise poses a major challenge in this domain, causing existing methods to suffer from two key limitations. First, they struggle to handle heterogeneous data noise, hindering robust multimodal representation learning. Second, they exhibit limited adaptability and generalization when encountering previously unseen noise. To address these issues, we propose Test-time Adaptive Hierarchical Co-enhanced Denoising Network (TAHCD). On one hand, TAHCD introduces the Adaptive Stable Subspace Alignment and Sample-Adaptive Confidence Alignment to reliably remove heterogeneous noise. They account for noise at both global and instance levels and enable jointly removal of modality-specific and cross-modality noise, achieving robust learning. On the other hand, TAHCD introduces Test-Time Cooperative Enhancement, which adaptively updates the model in response to input noise in a label-free manner, thus improving generalization. This is achieved by collaboratively enhancing the joint removal process of modality-specific and cross-modality noise across global and instance levels according to sample noise. Experiments on multiple benchmarks demonstrate that the proposed method achieves superior classification performance, robustness, and generalization compared with state-of-the-art reliable multimodal learning approaches.
keywords:
Multimodal classification, multimodal noise, multi-omics.
journal: Pattern Recognition
[…]
1 Introduction
With the rapid advancement of sensing and data processing technologies, large volumes of multimodal data have become available, spurring interest in multimodal learning and substantially improving performance across many applications [3, 21, 18], such as medical diagnosis [29, 40, 7] and cross-modality retrieval [20, 17, 27]. Data from different modalities provide complementary information that enables more comprehensive and discriminative representations for diverse tasks. However, they also inevitably introduce more complex noise, typically arising from sensor failure, environmental interference, and data collection errors. In real-world scenarios, multimodal noise exhibits two key characteristics: (1) Heterogeneity. Previous work [36] identifies two main types of multimodal noise: modality-specific noise and cross-modality noise. Modality-specific noise arises independently in each modality from sensor errors or environmental disturbances, while cross-modality noise stems from low-quality data due to weakly aligned or misaligned multimodal samples [6]. (2) Unpredictability. Noise often evolves over time as new samples are incrementally acquired, potentially introducing previously unseen noise. For example, unexpected device failures can introduce unfamiliar noise in newly collected patient data, image–text–mismatched posts continuously emerge in online data streams, and sensors in autonomous driving systems are exposed to varying weather and urban conditions.
@@ 1 Introduction
In the presence of complex multimodal noise, conventional multimodal methods struggle to produce reliable results, making them unsuitable for safety-critical applications such as disease diagnosis [13, 43, 41]. Accordingly, many methods improve the reliability of multimodal learning by removing noise and uninformative feature in data. Among them, [14, 13, 41, 43, 37, 4] address modality-specific noise via confidence-based feature learning and fusion. However, they generally assume well-aligned inputs and overlook the presence of cross-modality noise. Conversely, [15, 25, 34, 33] eliminate cross-modality noise via aligning and rectifying across modalities, but modality-specific noise can easily lead them to learn erroneous and unreliable alignments. Moreover, existing methods typically learn on limited training data, leaving them unaware of unseen noise. Therefore, these methods generally face the following limitations. (1) Insufficient robustness to heterogeneous noise: They struggle to jointly handle modality-specific and cross-modality noise. (2) Limited generalization: they adapt and generalize poorly to unseen noise.
@@
To address these limitations, we propose Test-time Adaptive Hierarchical Co-enhanced Denoising Network (TAHCD). TAHCD first removes heterogeneous noise by considering it at both global and instance levels. As shown in Figure 1, data often contain global-level noise with consistent patterns across samples, such as persistent sensor faults or repeated sampling under identical conditions, and instance-level noise affecting individual samples, e.g., sudden sensor failures or human errors. Accordingly, TAHCD proposes Adaptive Stable Subspace Alignment (ASSA) and Sample-Adaptive Confidence Alignment (SACA) to jointly eliminate modality-specific and cross-modality noise at the global and instance levels, respectively, achieving robust representation learning. Moreover, TAHCD incorporates Test-Time Cooperative Enhancement (TTCE) to collaboratively refine global- and instance-level denoising based on input noise in a label-free manner, enabling adaptation to previously unseen noise and substantially improving generalization.
@@
Specifically, ASSA first computes each modality’s feature covariance and performs singular value decomposition. A learnable mask conditioned on the singular values then selects eigenvectors to construct stable feature subspaces, guided by inter-class orthogonality and subspace projection alignment constraints. These constraints encourage the mask to discard directions carrying modality-specific and cross-modality noise across all samples. By aligning modality projections in the constructed stable subspaces, the subspace projection alignment in ASSA improves the reliability of cross-modality noise removal by preventing the model from erroneously aligning uninformative semantic content induced by modality-specific noise. Subsequently, SACA uses the globally denoised features from ASSA to estimate priors that guide the experts in removing heterogeneous noise at the instance level via a confidence-aware asymmetric slack alignment. By reweighting gradients based on modality confidence, this strategy encourages low-confidence modalities to align towards high-confidence ones, enabling reliable cross-modality noise removal under instance-level modality-specific noise. It also adopts a slack alignment scheme that constrains cross-modality discrepancies within a reasonable range learned from the prior, reflecting inherent differences in informative content across modalities. Compared with existing methods that strictly enforce cross-modality consistency, this prevents over-suppressing modality-specific complementary information while removing noise, resulting in more robust multimodal representations.
@@
Building upon the above framework for heterogeneous noise removal, we further propose TTCE that adapts the denoising framework to unseen noise in a label-free manner. Since the instance level can capture noise variations more sensitively than the global level, it can also perceive unseen noise more accurately. Based on this consideration, TTCE first incorporates the instance-level modality-specific and cross-modality noise learned by SACA into the reconstruction of globally denoised modality features produced by ASSA. This enables global-level modality representation learning in ASSA to explicitly leverage more accurate and finer-grained unseen-noise information captured at the instance level, thereby enhancing global-level unseen-noise removal. The enhanced global-denoised features in ASSA can provide more reliable priors, which in turn enhance instance-level unseen-noise learning and removal in SACA. Through multiple iterations of the above co-enhancement process between ASSA and SACA, the model achieves progressive update and adaptation to unseen noise.
@@
  • 3.
    The proposed TTCE in TAHCD cooperatively enhances ASSA and SACA driven by input sample noise in a label-free manner. This enables adaptive improvement of the denoising process at test time for previously unseen noise, thereby enhancing generalization ability.
2 Related Works
@@
2.1 Multimodal Learning
The advancement of sensor technology and the development of data transmission methods have driven the rapid growth of multimodal learning in a wide range of tasks and applications [3, 42]. For instance, [19, 29, 40, 7] have achieved precise pathological classification by integrating multiple omics data. [20, 17, 27] have developed cross-modality retrieval methods based on modalities such as images and text. [31, 38] have constructed multimodal networks based on images and text, improving performance on computer vision tasks such as classification and generation. However, in real-world scenarios, the prevalence of low-quality data caused by noise substantially undermines the reliability of multimodal methods [36, 37], with particularly severe consequences in safety-critical applications such as medical diagnosis. Therefore, this work focuses on low-quality data and develops multimodal learning method that is robust and generalizable to multimodal noise.
2.2 Reliable Learning from Low-Quality Multimodal Data
@@
In real-world scenarios, multimodal learning often suffers from low-quality data, which undermines reliability. A major source of low-quality data is multimodal noise, typically arising from factors such as sensor failures, environmental influences, or human errors. Existing studies [36] categorize multimodal noise into two types, namely modality-specific noise and cross-modality noise, and have developed various methods to address these noise types for reliable multimodal learning.
Many studies mitigate modality-specific noise through quality-based representation learning and fusion methods. Geng et al. [12] have proposed the DUA-Nets, which achieved uncertainty-based multimodal representation learning through reconstruction. Han et al. [13] have modeled informativeness at the feature and modality levels, achieving trustworthy multimodal feature fusion. Zhang et al. [37] have achieved more robust multimodal fusion by dynamically assigning weights to each modality based on uncertainty estimation. Zheng et al. [39] have achieved trustworthy multimodal classification via integrating feature and label-level confidence. Zhou et al. [41] have introduced a trustworthy multi-view classification framework via confidence-aware fusion. Cao et al. [4] have proposed adjusting modality weights based on their losses to achieve robust fusion. However, these methods assume that the modalities are well-aligned and therefore tend to yield unreliable results in the presence of cross-modality noise. Conversely, many studies delve into cross-modality noise removal by modeling cross-modality alignment and correspondence. [26, 11] have proposed rule-based methods to remove misaligned or weakly aligned data. [16, 8, 33, 32] have designed models to filter or correct misaligned samples. [15, 34, 9] have proposed regularization methods to mitigate the impact of misaligned samples. However, these methods are easily misled by modality-specific noise, leading to erroneous and unreliable alignment learning. Moreover, existing methods typically learn and update on limited training data, hindering adaptation and generalization to unseen and varying noise conditions. Therefore, we propose a multimodal learning method that jointly handles modality-specific and cross-modality noise and enables test-time adaptive updates, thus enhancing robustness to heterogeneous noise and generalization to unseen noise.
3 Proposed Approach
@@
3.1 Overview of TAHCD
As illustrated in Figure 2, TAHCD first employs modality-specific feature encoders to extract representations from each modality. It then introduces Adaptive Stable Subspace Alignment (ASSA) to remove modality-specific and cross-modality noise exhibiting global distributional characteristics. In addition, Sample-Adaptive Confidence Alignment (SACA) is proposed to eliminate modality-specific and cross-modality noise at the instance level. Building upon these components, we further design Test-Time Cooperative Enhancement (TTCE) mechanism. It achieves cooperative enhancement between global- and instance-level denoising, driven by sample noise in a label-free manner, improving adaptability and generalization to unseen noise.
3.2 Adaptive Stable Subspace Alignment
@@
3.2.1 Adaptive Stable Subspace Construction
To facilitate global-level multimodal noise removal, we adaptively eliminate statistically unstable and uninformative directions in the feature distribution of each modality and suppress noise-dominated variations across samples. Specifically, for the $m$-th modality, the input $x^{m}=\{x_{1}^{m},\dots,x_{N}^{m}\}$ is first encoded into the latent space $z^{m}=\{z_{1}^{m},\dots,z_{N}^{m}\}$ by the corresponding modality encoder $\phi_{x}^{m}$: $z^{m}=\phi_{x}^{m}(x^{m})$. Then, the covariance matrix $\Sigma^{m}_{z}$ among features is computed from $z^{m}$, which can be formulated as:
@@ 3.2.1 Adaptive Stable Subspace Construction
$\phi_{\lambda}^{m}(\cdot)$ is a learnable encoder, $\sigma(\cdot)$ denotes the sigmoid function. $h^{m}=\{h_{1}^{m},\dots,h_{N}^{m}\}$ denote the globally denoised representation obtained from the latent feature $z^{m}$ of modality $m$. As shown in Eq. (5), the feature vector $z^{m}$ is first projected onto the covariance principal axes using the orthogonal basis $U_{z}^{m}$. It is then filtered by the learned mask $w^{m}_{\lambda}$ and finally mapped back to the original dimension via $(U^{m}_{z})^{\top}$.
3.2.2 Constraints
To ensure that the adaptive stable subspace construction sufficiently eliminates modality-specific and cross-modality noise, we introduce inter-class orthogonality constraint $\mathcal{L}_{\mathrm{o}}$ and subspace projection alignment constraint $\mathcal{L}_{\mathrm{a}}$. The inter-class orthogonality constraint $\mathcal{L}_{\mathrm{o}}$ eliminates modality-specific noise by removing the shared spurious patterns it induces across classes within the same modality, while preserving discriminative features of different classes, which can be formulated as:
@@ 3.2.2 Constraints
3.3 Sample-Adaptive Confidence Alignment
To provide reliable supervision signals for instance-level noise removal, we first estimate modality-wise feature distributions and cross-modality discrepancy distributions using the globally denoised representations obtained from ASSA. These distributions are then serve as priors to guide sample-adaptive noise experts with confidence-aware asymmetric slack alignment for instance-level noise removal.
3.3.1 Prior Estimation